HW: Chapter 5 (5.3 & 5.4) i Saved Find the EAR in each of the following cases. (Do not round intermediate calculations and enFind the EAR in each of the following cases. (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16. Use 365 days in a year.)
Answer #1
The formula for Effective annual interest rate (EAR) is
EAR = (1+i/n)^n - 1
where i is the stated rate
and n is the number of compounding periods
1)
i = 9.75%
n = 4
EAR = (1+.0975/4)^4 - 1
EAR = (1+.024375)^4 - 1
EAR = (1.024375)^4 - 1
EAR = 1.101123 - 1
EAR = .101123
The EAR is 10.11%.
2)
i = 16.25%
n = 12
EAR = (1+.1625/12)^12 - 1
EAR = (1.013542)^12 - 1
EAR = (1.175166 - 1)
EAR = .175166
The EAR is 17.52%.
3)
i = 15.75%
n = 365
EAR = (1+.1575/365)^365 - 1
EAR = (1+.000432)^365 - 1
EAR = (1.000432)^365 - 1
EAR = 1.170541 - 1
EAR = .170541
The EAR is 17.05%.
4)
i = 11.75%
n = 2
EAR = (1+.1175/2)^2 - 1
EAR = (1.05875)^2 - 1
EAR = 1.120952 - 1
EAR = .120952.
The EAR is 12.09%.
Answer #2



SOLUTION :


Stated rate                 Number of times                   Effective rate

  (APR)                          compounded                            (EAR)



  9.75%                              Quarterly                  (1 + 0.0975/4)^4 - 1 = 0.1011 = 10.11%


 16.25%                             Monthly                 (1 + 0.1625/12)^12 - 1 = 0.1752 = 17.52%


 15.75%                             Daily                  (1 + 0.1575/365)^365 - 1 = 0.1705 = 17.05%


 11.75%                     semiannually                     (1 + 0.1175/2)^2 - 1 = 0.1210 = 12.10% 

answered by: Tulsiram Garg